What is the temperature If the peak of a blackbody
Answer #1

2. Given

   the energy required to break hydrogen bond in protein molecule is

   E = 0.1 eV
we know that E = h*f ==> f = E/h
we know that 1 eV = 1.6*!0^-19 J, and h is planck's constant = 6.626*10^-34 m2 kg/s

       f = (0.1*1.6*10^-19) /(6.626*10^-34) = 2.4147298520978*10^13 Hz


the wavelength is L = c/f c is speed of light

       L = (3*10^8)/(2.41473*10^13) m

       L = 1.242375*10^-5 m

the maximum wavelength is 124233 nm